pls advise how to solve this:
a ladder 10 ft. long is leaning against a building. how high on the building will the ladder be when the bottom of the ladder is 6ft from the building?








Answer by stanbon(75887)

a ladder 10 ft. long is leaning against a building. how high on the building will the ladder be when the bottom of the ladder is 6ft from the building?
------------------------------
the building, the ladder and the ground form a right triangle with
the ladder being the hypotenuse and the ground being 6 ft.
Draw the rectangle
Let the height of the building from the ground to the top of the
ladder be "h".
Use Pythagoras to get:
h^2 + 6^2 = 10^2
h^2 = 100-36 = 64
h = 8 ft
-----------
Cheers,
Stan H.